OpenStudy (anonymous):

-9+x=x/8-23 two step equation...solve for x

OpenStudy (xxdragoonxx):

-9+x=x/8-23 LCM 8 x-184=8x-72 72-8x+72=x-184+72 x=8x-112 112-x=x-8x -112=7x -112/7=7x/7 -16=x